Jump to main content or area navigation.

Contact Us

Waste Site Cleanup & Reuse in New England

Atlas Tack Corporation Superfund Site
Bi-Weekly Remedial Activities Update
Report # 21: December 4 - December 15, 2006

« Back to Atlas Tack Superfund Site Fact Sheet

1. ACTIVITIES PERFORMED DURING THE WEEKS ENDING 8 DEC. – 15 DEC.

  • Removed all of the Subtitle D soils and oversize debris generated to date and transported to the TREE landfill in Rochester, NH
  • Continue transport and disposal (T&D) of COMM-97- soils to Marion from stockpiles
  • Completed the excavation of soils in the Northwest Fill Area, with the exception of the 1,500-cy of soils remaining in the sliver left in place to prevent Boy’s creek from infiltrating the excavation areas
  • Constructed temporary haul road though the Lagoon Area
  • Began to excavate soils along the eastern cell of the Lagoon Area, removed the top 4-ft along part of the area
  • Began to collect stream gauge flow measurements in Boys Creek; collected data will be used to determine dilution factor for surface water discharge
  • Collect baseline water samples from (3) locations within Boys Creek, (1) from standing water in the infiltration basin and (1) from standing water in the Commercial and Industrial Debris Area (CID)
  • Collected samples to the fence-line along the western boundary of the

Ongoing Site Activities:

  • Collected waste characterization soil samples from soils originating from NW-Fill and Lagoon areas and Lagoon for disposal profiling for off site landfill disposal
  • Collected confirmatory soil samples from the floors of the NW-Fill Area
  • Continue water handling activities
  • Continued to inspect trucks prior to exiting the site
  • Continued to provide flaggers to cross trucks across the bike paths along Adams and Tripp Streets
  • Continued stockpile management activities
  • Continue with air monitoring program

2. PROJECTED ACTIVITIES FOR THE WEEKS ENDING 22 DEC. – 29 DEC.

  • Observe December 26th as Christmas Holiday
  • Continue with excavation activities in the Lagoon Area
  • Begin to backfill the eastern cell of the Lagoon Area
  • Continued to T&D COMM-97 and Subtitle D soils
  • Collect waste characterization soil samples from soils originating from Lagoon Area

Ongoing Site Activities:

  • Continue to collect pre and post excavation survey data of the excavations
  • Continue with stockpile management activities
  • Continue water handling activities
  • Continued to inspect trucks prior to exiting the site
  • Continued to provide flaggers to cross trucks across the bike paths along Adams and Tripp Streets
  • Continue with air monitoring program

3. AIR MONITORING INFORMATION

WESTON monitored for particulate (dust) during intrusive activities performed during this reporting period. WESTON recognizes EPA's level of concern for "PM10" to caution sensitive groups, including children, the elderly, and people with heart or lung disease, about exposure to coarse dust particles. The PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure.

Perimeter Air Monitoring is performed at up to (6) locations throughout a given work day. Readings are collected at 60 second intervals throughout the day and reported as a daily average.

If exceedances are identified, an increased effort of engineering controls (water suppression) will be utilized to adhere to the PM10 level of concern. If engineering controls appear to be inadequate, any dust generating activities will stop. Current dust control practices will be re-evaluated until concentrations are reduced to below the PM10 standard.

Results for 30 November – 13 December 2006 can be reviewed in the attached table.

Weekly Dust Monitoring Summary
Week #21

The EPA's PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure

Date Pleasant Street (µg/m3) Adams Street (µg/m3) Tripp Street (µg/m3) Church Street (µg/m3)
11/30/2006 0 0 5 0
12/1/2006 Precip Precip Precip Precip
12/4/2006 Precip Precip Precip Precip
12/5/2006 0 0 0 0
12/6/2006 0 0 0 0
12/7/2006 0 0 0 0
12/8/2006 Precip Precip Precip Precip
12/11/2006 0 0 0 0
12/12/2006 0 0 0 0
12/13/2006 0* 0* 6* 0*

Notes:
All readings reflect the daily average calculated over the course of the workday.
All readings collected with a Thermo Electron PDR 1000.
NSC: No sampling conducted; No intrusive work conducted in that area.
Precip: No sampling conducted due to precipitation.
*: Readings are for 4hrs as PDR's were taken down due to precipitation

4. T&D "BI-WEEKLY" SUMMARY

DATE
DESTINATION
WASTE DETERMINATION
TONS
CY
STOCKPILES
13 Dec. 06 MARION COMM-97 640.00 426.67 WF-07, 09, 10
13 Dec. 06 TREE Subtitle-D/OVERSIZE 360.00 240.00 EA-20, 28, 37
12 Dec. 06 TREE Subtitle-D/OVERSIZE 377.31 251.54 EA-20, 28, 37
11 Dec. 06 TREE Subtitle-D/OVERSIZE 394.3 262.87 EA-20, 28, 37
06 Dec. 06 MARION COMM-97 428.50 285.67 CD-17, 18, 20, 22; EA-16, 17, 19
06 Dec. 06 TREE Subtitle-D 441.11 294.07 CD-17, 18, 20, 22; EA-16, 17, 19
05 Dec. 06 MARION COMM-97 715.08 476.72 CD-25, WF-03
05 Dec. 06 TREE Subtitle-D 533.48 355.65 EA-30, 32, 35, 06
04 Dec. 06 TREE Subtitle-D 831.22 554.15 EA-30, 32, 35
01 Dec. 06 TREE Subtitle-D 1059.95 706.63 WF-05
30 Nov. 06 TREE Subtitle-D 921.22 614.15 CD-17, 18, 20, 22; EA-16, 17, 19
29 Nov. 06 TREE Subtitle-D 627.84 418.56 CD-17, 18, 20, 22; EA-16, 17, 19
28 Nov. 06 TREE Subtitle-D 786.3 524.20 CD-17, 18, 20, 22; EA-16, 17, 19
27 Nov. 06 TREE Subtitle-D 650.66 433.77 CD-17, 18, 20, 22; EA-16, 17, 19

5. T&D TOTAL SUMMARY

WK/E
DESTINATION
WASTE DETERMINATION
ORIGIN
CY
TONS
*15 Dec. 06 TREE Sub-D Oversize EAST FILL 693.00 1,040.00
*15 Dec. 06 Marion COMM-97 NW-Fill 756.00 1,134.00
8 Dec. 06 TREE Sub-D Soils EAST FILL 1,203.87 1,805.81
8 Dec. 06 Marion COMM-97 NW-Fill/CID 762.39 1,143.58
1 Dec. 06 TREE Sub-D Soils CID/EAST/NW FILL 2,697.13 4,045.97
24 Nov. 06 TREE Sub-D Soils CID/ EAST FILL 1,060.00 1,590.00
17 Nov. 06 Bourne Sub-D Oversize CID/ EAST FILL 1,981.00 2,481.00
10 Nov. 06 TREE SUB-D CID 1,130.37 1,695.56
03 Nov. 06 Marion COMM-97 CID/EAST FILL 1,881.51 2,822.27
27 Oct. 06 T&D DID NOT OCCUR THIS WEEK
20 Oct. 06 T&D DID NOT OCCUR THIS WEEK
13 Oct. 06 T&D DID NOT OCCUR THIS WEEK
06 Oct. 06 Marion COMM-97 CID 843.04 1,264.56
29 Sept. 06 Marion COMM-97 CID 449.63 674.45
22 Sept. 06 Crapo COMM-97 CID 291.23 436.85
15 Sept. 06 Crapo COMM-97 CID 533.67 800.51
T&D TOTAL
14,283.17 20,934.56

Notes:
* Total CY and Tons are based on estimates and will be updated at the end of the week

6. EXCAVATION SUMMARY

WEEKS 12/08/06 – 12/15/06
TO DATE
ACTIVITIES INITIAL TOTAL QUANTITY INITIAL TOTAL QUANTITY
Commercial Area Excavation Complete 2,655 CY
C.I.D. Excavation None 7,083 CY
East Fill Excavation Complete 9,046 CY
NW-Fill Area 812 CY 8,612 CY
Lagoon Area 2,000 CY 2,000 CY
 

Jump to main content.